OpenStudy (anonymous):

NEED ANSWER ! Select the inequality of the graph shown below. graph of an absolute value inequality. There are points at (-12, 1), (-11, 2), (-10, 3), (-9, 2), and (-8, 1). Shading is below the solid line Answer y less than or equal to – |x + 10| + 3 y less than or equal to – |x – 10| + 3 y greater than or equal to – |x – 10| + 3 y greater than or equal to – |x + 10| + 3

OpenStudy (anonymous):

?

OpenStudy (ash2326):

wait , I'm solving

OpenStudy (anonymous):

sorry

OpenStudy (ash2326):

no problem

OpenStudy (ash2326):

there are two equations y=-|x-10|+3 and y=-|x+10|+3 we have to find which one satisfies the point given let's take -12,1 put x=-12 in both check what's the value of y

OpenStudy (anonymous):

okay

OpenStudy (ash2326):

y=-|-12-10|+3 y=-25 so y=-|x-10|+3 is not our equation let's check the other one y=-|x+10|+3 y=-|-12+10|+3 y=1 so y=-|x+10|+3 is the required so either it's option first or last, we'll check that

OpenStudy (anonymous):

so is it the first one ?

OpenStudy (ash2326):

we'll check, sorry my internet got disconnected

OpenStudy (ash2326):

when x<-10 y=x+13 when x>=-10 y=-x-7 let's plot these

OpenStudy (ash2326):

 now let's take a point 4, 4 which is above put this in y=-x-7 4>=-11 that is above the line, we are given shade region below the line so y<=-|x+10|+3 option 1

OpenStudy (ash2326):

sorry, I kept you waiting:c

OpenStudy (mertsj):